# Understanding and Improving Shampoo and SOAP via Kullback–Leibler Minimization

### v3

@@ 6 Experimental Setup and Empirical Evaluations
Figure 5: Empirical results—based on random search with 150 runs per method—demonstrate the advantages of KL-Shampoo’s (two-sided) estimation over other Shampoo variants under comparable settings for NN training, including Shampoo with $p=\nicefrac{{1}}{{2}}$ (one-sided, no grafting), F-Shampoo (two-sided, Frobenius-norm–based), and VN-Shampoo (trace scaling, two-sided von-Neumann-divergence-based). We make these variants practical by incorporating a QR step and an EMA scheme for eigenvalue estimation (Fig.˜3). To ensure a fair comparison and minimize implementation bias, we implement Shampoo, F-Shampoo, and VN-Shampoo ourselves, aligning them closely with KL-Shampoo. See Fig.˜10 (Appx.˜H) for a detailed comparison between KL-Shampoo and VN-Shampoo.
@@
In the last set of experiments, we evaluate the benefits of using the two-sided estimation scheme under our KL perspective. Specifically, we compare the two-sided approach (KL-Shampoo) against the one-sided approach (Shampoo) in a comparable setting. To ensure fairness and eliminate implementation bias, we use our own implementation of Shampoo aligned closely with that of KL-Shampoo. For this comparison, we extend Shampoo with a QR-based step and our EMA scheme for eigenvalue estimation, as described in Fig.˜3. As shown in Fig.˜5, KL-Shampoo consistently and significantly outperforms Shampoo, even when Shampoo employs a similar QR-based estimation rule.
@@ Appendix C Proof of ˜3
Because the weight matrix is block-diagonal, we can slice this objective function for the proximal step into two terms.
@@
$\quad=\underbrace{\langle\nabla_{{\bm{S}}_{a}}\mathcal{L}\big|_{{\bm{S}}={\bm{S}}^{(t)}},\,{\bm{X}}_{a}\rangle+\frac{1}{2\beta_{2}}\|{\bm{X}}_{a}-{\bm{S}}_{a}^{(t)}\|_{{\bm{W}}_{a}}^{2}}_{\text{(block ${\bm{X}}_{a}$)}}+\underbrace{\langle\nabla_{{\bm{S}}_{a}}\mathcal{L}\big|_{{\bm{S}}={\bm{S}}^{(t)}},\,{\bm{X}}_{b}\rangle+\frac{1}{2\beta_{2}}\|{\bm{X}}_{b}-{\bm{S}}_{b}^{(t)}\|_{{\bm{W}}_{b}}^{2}}_{\text{(block ${\bm{X}}_{b}$)}}$
@@ Appendix G Key Distinction between Shampoo with trace scaling and KL-Shampoo
We now introduce the definition of a Bregman divergence (Bregman, 1967) to formally discuss the distinction between Shampoo with trace scaling and KL-Shampoo. Given a strictly convex and differentiable (scalar) function $F(\cdot)$, the Bregman divergence based on this function is defined as
@@ Appendix H Additional Experiments
We conduct three additional sets of experiments, following the same experimental setup as described in the main text, to further evaluate our approach. Due to our limited computation, we focus on two language models: NanoGPT (123M) and Llama (134M) in these additional experiments.

### v4

@@ 6 Experimental Setup and Empirical Evaluations
Figure 5: Empirical results—based on random search with 150 runs per method—demonstrate the advantages of KL-Shampoo’s (two-sided) estimation over other Shampoo variants under comparable settings for NN training, including Shampoo with $p=\nicefrac{{1}}{{2}}$ (no grafting, Eq.˜1), F-Shampoo (two-sided, Frobenius-norm–based, Fig.˜6), and VN-Shampoo (trace scaling, two-sided von-Neumann-divergence-based, Fig.˜8). We make these variants practical by incorporating a QR step and an EMA scheme for eigenvalue estimation (Fig.˜3). To ensure a fair comparison and minimize implementation bias, we implement Shampoo, F-Shampoo, and VN-Shampoo ourselves, aligning them closely with KL-Shampoo. See Fig.˜10 (Appx.˜H) for a detailed comparison between KL-Shampoo and VN-Shampoo.
@@
In the last set of experiments, we evaluate the benefits of using the two-sided estimation scheme under our KL perspective. Specifically, we compare the two-sided approach (KL-Shampoo) against the one-sided approach (Shampoo) in a comparable setting. To ensure fairness and eliminate implementation bias, we use our own implementation of Shampoo aligned closely with that of KL-Shampoo. For this comparison, we extend Shampoo with a QR-based step and our EMA scheme for eigenvalue estimation, as described in Fig.˜3. Similarly, we also consider two more Shampoo variants discussed in Sec.˜3.3. As shown in Fig.˜5, KL-Shampoo consistently outperforms other Shampoo variants, even when these variants employ a similar QR-based estimation rule.
@@ Appendix C Proof of ˜3
Because the weight matrix $\bm{W}$} is block-diagonal, we can slice this objective function for the proximal step into two terms.
@@
$\quad=\underbrace{\langle\nabla_{{\bm{S}}_{a}}\mathcal{L}\big|_{{\bm{S}}={\bm{S}}^{(t)}},\,{\bm{X}}_{a}\rangle+\frac{1}{2\beta_{2}}\|{\bm{X}}_{a}-{\bm{S}}_{a}^{(t)}\|_{{\bm{W}}_{a}}^{2}}_{\text{(block ${\bm{X}}_{a}$)}}+\underbrace{\langle\nabla_{{\bm{S}}_{b}}\mathcal{L}\big|_{{\bm{S}}={\bm{S}}^{(t)}},\,{\bm{X}}_{b}\rangle+\frac{1}{2\beta_{2}}\|{\bm{X}}_{b}-{\bm{S}}_{b}^{(t)}\|_{{\bm{W}}_{b}}^{2}}_{\text{(block ${\bm{X}}_{b}$)}}$
@@ Appendix G Key Distinction between Shampoo with trace scaling and KL-Shampoo
We now introduce the definition of a Bregman divergence to formally discuss the distinction between Shampoo with trace scaling and KL-Shampoo. Given a strictly convex and differentiable (scalar) function $F(\cdot)$, the Bregman divergence based on this function is defined as
@@ Appendix H Additional Experiments
We conduct three additional sets of experiments, following the same experimental setup as described in the main text, to further evaluate our approach. Due to limited computational resources, we focus on two language models—NanoGPT (123M) and Llama (134M)—in these additional experiments.